OpenStudy (anonymous):

Can someone explain to me how to do this lol, i'm kinda slow. Solve: −3x−3 =12

OpenStudy (anonymous):

start by adding \(3\) to both sides, to get rid of the \(-3\) on the left

OpenStudy (anonymous):

add 3 to both sides, it will get rid of it on the side it is originating from, then divide by -3 and you will get -5

OpenStudy (anonymous):

you get \[-3x-3=12\] add 3 to get \[-3x=12+3\] or \(-3x=15\)

OpenStudy (anonymous):

First you would add three to both sides. -3x = 15 then divide by -3 x = -5

OpenStudy (anonymous):

once you have \(-3x=15\) divide both sides by \(-3\) do get \[x=15\div -3\] or \(x=-5\)

OpenStudy (anonymous):

Thank you all for the help :)

OpenStudy (anonymous):

no problem